Maxx Industry is regularly supplying some critical parts to your assembly line. These parts are either good or defective. During each delivery a sample of 500 units is sampled and the number of defective units in the sample is noted. The average defect rate is 3.5%.

a. Develop the upper and lower control limits (z=3).

b. What is process capability index and what will be your objective in finding the process capability?



1
Expert's answer
2021-11-22T17:03:02-0500

Answer

a) Upper lower limit

Standard deviation (z) = 3

Average = 3.5% of 500 which is 17.5

Multiply standard deviation by 3, which is 9,

Add the multiplied number to the average to get the upper-class limit,

     "9+17.5 = 26.5" ,

Therefore upper-class limit is 26.5 Units.

Lower Class Limit

Multiply standard deviation by 3

"3*3=9"

Subtract nine from the average, which is 17.5

  "17.5 - 9= 8.5"

Therefore the lower class limit is 8.5 units.

Process capability index is a set of calculations that determines whether, statistically, a system is able to meet a set of specification limits.

This process capability analysis will help ensure that the parts from Maxx Industry align with what my assembly requires.